# Entangled in Representations: Mechanistic Investigation of Cultural Biases in Large Language Models

### v1

@@ Abstract
The growing deployment of large language models (LLMs) across diverse cultural contexts necessitates a better understanding of how the overgeneralization of less documented cultures within LLMs’ representations impacts their cultural understanding. Prior work only performs extrinsic evaluation of LLMs’ cultural competence, without accounting for how LLMs’ internal mechanisms lead to cultural (mis)representation. To bridge this gap, we propose Culturescope, the first mechanistic interpretability-based method that probes the internal representations of LLMs to elicit the underlying cultural knowledge space. CultureScope utilizes a patching method to extract the cultural knowledge. We introduce a cultural flattening score as a measure of the intrinsic cultural biases. Additionally, we study how LLMs internalize Western-dominance bias and cultural flattening, which allows us to trace how cultural biases emerge within LLMs. Our experimental results reveal that LLMs encode Western-dominance bias and cultural flattening in their cultural knowledge space. We find that low-resource cultures are less susceptible to cultural biases, likely due to their limited training resources. Our work provides a foundation for future research on mitigating cultural biases and enhancing LLMs’ cultural understanding. Our codes and data used for experiments are publicly available [^1] https://github.com/copenlu/CultureScope.
@@ 1 Introduction
Large language models (LLMs) are increasingly being used in culturally diverse contexts, where understanding and responding appropriately to each culture is essential (Pandya and Holia, 2023; Salemi et al., 2023; Liu et al., 2025). However, the cultural knowledge acquired by LLMs is largely shaped by the data they are trained on, which is predominantly Western-centric (Santurkar et al., 2023). This results in severe geographic imbalances, where some regions receive disproportionate attention while others remain marginalized (Nguyen et al., 2023a). While LLMs do acquire some cross-cultural patterns during training (Hershcovich et al., 2022; Arora et al., 2023), this learning often results in overgeneralization. For example, when asked about a popular leisure activity for retired men in Azerbaijan, GPT-4 responded with chess—a plausible answer, but one that reflects general stereotypes of post-Soviet or Eastern European regions rather than Azerbaijan specifically (Myung et al., 2024).
@@
Figure 1: Given the question about the popular domestic vacation spot in Greece, Culturescope first generates an answer to cultural Question at the Inference stage. Then, it reads the hidden representation from the Inference stage and elicits the cultural knowledge used for ‘Zakynthos’ at the Scoping-in stage. We finalize a list of cultural knowledge after the Filtering stage. Culturescope unveils the internal mechanism of LLMs that cannot be revealed through the Inference stage alone.
@@
However, these patterns of overgeneralization directly perpetuate cultural biases such as cultural flattening. Measuring such overgeneralizations through model outputs alone proves insufficient since generated outputs mask the underlying mechanisms that lead to cultural (mis)representation. To facilitate the examination of the underlying mechanisms, we propose to study LLMs’ cultural understanding with mechanistic interpretability (MI) techniques. MI techniques provide us with methods that can directly examine how cultural biases discovered by the extrinsic evaluation Santurkar et al. (2023) is internally processed within model representations, revealing where and how harmful generalizations emerge. We are the first to propose an approach for intrinsic cultural bias evaluation.
@@
In this work, we introduce Culturescope, a method to probe internal representations and surface the cultural knowledge activated during cultural understanding tasks (§4.1). By surfacing this knowledge, we reveal the internal cultural knowledge space from which model outputs are generated. Figure 1 illustrates an overview of Culturescope. To examine the intrinsic ‘cultural flattening’ embedded in the parameter space, we introduce a cultural flattening (CF) score, which quantifies the degree of intersection between cultural knowledge decoded by Culturescope (§4.2). We implement our framework on two cultural understanding tasks, cultural commonsense Question Answering (QA) and extractive QA, across three different LLMs.
@@
We further challenge the model’s cultural understanding by creating multiple-choice questions (MCQs) with hard negatives (§3.2.3). Culturally nuanced answers from high-resource cultures or geographically proximate countries are selected as hard negative options to simulate the Western-dominance bias and the cultural flattening. This setup refrains LLMs to leverage surface-level elimination strategies based on the overgeneralization Khan et al. (2025). Analyzing selected options by LLMs with the attention map method Yuksekgonul et al. (2024) allows us to examine whether extrinsic and intrinsic cultural biases align, by revealing which options the model internally attends to (§5.1).
High CF scores of Western and high-resource cultures show that LLMs encode Western-dominance bias and cultural flattening in their cultural knowledge space (§6.2). Model performances on MCQs with hard negatives demonstrate that low-resource cultures are less likely to be affected by the biases, likely due to their limited training resource (§6.3). This implies that LLMs struggle with low-resource cultures due to the lack of parametric knowledge. In §6.4, we find that LLMs over-attend to tokens from Western and high-resource cultures. This suggests Western-dominance bias is more internalized than the cultural flattening. Our work provides useful signal to future research on mitigating the internalized patterns for biases to build a better culturally aligned model.
@@ 2 Related Work
Previous work has proposed evaluation datasets and frameworks to assess LLMs’ cultural understanding ability acquired during pre-training Keleg and Magdy (2023); Naous and Xu (2025); Pawar et al. (2025). BLEnD Myung et al. (2024) provides a multilingual commonsense QA dataset spanning 16 countries and regions, designed to uncover cross-cultural disparities in everyday knowledge. CAMeL Naous et al. (2024) compares LLM behavior in Arabic versus Western settings across tasks like story generation, NER, and sentiment analysis, exposing systematic cultural biases in LLMs. Other multilingual benchmarks Zhou et al. (2025); Hasan et al. (2025); Wang et al. (2024); Cao et al. (2024) construct culturally localized evaluation datasets that span domains such as cuisine, proverbs, news, and reasoning. Across these datasets, performance gaps are consistently observed between high-resource and underrepresented languages and cultures, often linked to pre-training data imbalances that favor dominant regions Naous and Xu (2025).
@@
MI techniques are developed to explain the inner workings of LLMs by identifying responsible model components, such as neurons and attention heads Meng et al. (2022); Geva et al. (2023); Yu et al. (2024). Leveraging their transparency, recent studies have employed MI techniques to investigate how specific behaviors emerge in LLMs. For instance, they have been used to uncover and manipulate components associated with social biases, enabling both diagnostic and steering interventions Liu et al. (2024); Durmus et al. (2024); Yang et al. (2024). Despite growing interest in the cultural capabilities of LLMs, no prior work has explored cultural understanding through the lens of MI. Our study fills this gap by applying MI techniques to probe the internal representation of cultural knowledge in LLMs, offering new insights into how cultural understanding is encoded and organized within the model.
@@ 3.1 Preliminaries
A dataset $D=[(q_{1},C_{1},a_{1}),...,(q_{N},C_{N},a_{N})]$ consists of $N$ tuple instances containing: a question $q$, an option list $C=[c_{0},c_{1},c_{2},a]$ containing three options for MCQ, one gold answer $a$ and a country of interest $y$. For MCQ, an LLM is given $C$ and $q$ to generate an output $O=[o_{0},...,o_{P}]$ consisting of $P$ tokens. For open-ended QA, an LLM is only given $q$ to generate an output $O$.
To generate an answer, an LLM converts a tokenized input text $T=[t_{0},...,t_{S}]$ containing $S$ number of tokens into $d$-dimenstional vectors using the embedding matrix $E\in\mathbb{R}^{|\mathbb{V}\times d|}$. Then, the vectors are processed through $L$ layers, each containing a multi-head self-attention (MHSA) layer and an MLP layer. The hidden representation $x_{i}^{l}$ from a layer $l$, on a token $t_{i}$ is computed by:
@@ 3.2.1 Cultural QA Datasets
We select BLEnD Myung et al. (2024), a cultural commonsense QA dataset, and CAMeL-2 Naous and Xu (2025), an extractive QA dataset featuring culturally grounded entities. BLEnD Myung et al. (2024) is a hand-crafted benchmark designed to evaluate LLMs’ everyday knowledge across diverse cultures in English. It comprises 500 short-answer question-answer pairs for each country, where the answers vary depending on the country’s cultural or regional context. To reduce computational cost, we exclude North Korea and West Java, resulting in a final selection of 14 countries from the BLEnD dataset.
CAMeL-2 Naous and Xu (2025) is a bilingual benchmark originally constructed to evaluate LLMs’ entity extraction capabilities on Arabic and English entities. With CAMeL-2, an LLM is asked to extract an entity from a context collected from Arabic X/Twitter data according to the specified entity type in the input. We take the English partition and reduce the dataset to 14 countries to keep a similar country distribution to the BLEnD dataset. Dataset details, including domains and countries covered by the datasets, can be found in Appendix A.
@@ 3.2.2 Grouping of Cultures
We categorize 14 countries from each dataset along the resource dimension and the region dimension to study how overgeneralization manifests across these dimensions. For the resource dimension, we adopt the taxonomy proposed by Joshi et al. (2020), which classifies languages into six levels (0: very low-resource to 5: very high-resource). For our experiments, we simplify this into three groups: High (Level 5), Mid (Levels 3-4), and Low (Levels 0-2). We assign each country a language resource level based on its most widely spoken language provided by Wikipedia. For the region dimension, we group countries into six regions based on continents. We split Asia into three subregions, which leaves us six regional groups: North America, Europe, Africa, West Asia, South Asia, and East Asia. A complete list of countries within each group is provided in Appendix A.1.
@@ 3.2.3 Cultural MCQ with hard negatives
Khan et al. (2025) found that if MCQs lack the adversarial depth to probe genuine cultural understanding, models can exploit surface-level elimination strategies without truly understanding cultural distinctions. Thus, we propose a cultural MCQ with hard negatives to study how overgeneralization—driven by regional or resource dominance or similarity—affects the downstream task. Since BLEnD Myung et al. (2024) provides different answers from each culture with the same question, we create BLEnD-resource and BLEnD-region partition using culturally nuanced answers in BLEnD.
@@
We design two types of multiple choice question option lists that incorporate hard negative options: $C_{resource}$ and $C_{region}$, corresponding to BLEnD-resource and BLEnD-region, respectively. For $C_{resource}$, given a question $q$ targeting country $y$, we sample one country from each of the three resource levels excluding $y$. We obtain these three country’s respective gold answers when substituted into $q$ for $y$, resulting in three hard negative options: $c_{high}$, $c_{mid}$, and $c_{low}$. For $C_{region}$, we sample one country from the same geographical region as $y$ (excluding $y$) and extract its corresponding answer to construct a region-based hard negative option, $c_{flatten}$. Two additional options are randomly selected from countries in different regions. All options are shuffled to avoid positional bias. Figure 2 shows the example of $C_{resource}$ and $C_{region}$ for the question about China’s popular family game. Cultural MCQ with hard negative options allows us to examine when a model generates incorrect answers to cultural questions, whether a model’s cultural confusion arises from similarity in resource level or regional proximity.
@@ 3.3 Models
Application of MI methods requires a full-access to the model weights. Due to the requirement, we conduct our experiments with three recent open-sourced LLMs: Meta-Llama-3.1-8B-Instruct (Llama-3.1, Grattafiori et al. (2024)), aya-expanse-8b (aya-expanse, Dang et al. (2024)), and Qwen2.5-7B-Instruct (Qwen2.5, Qwen Team (2024)).
@@ 3.4 Patchscope
Patchscope Ghandeharioun et al. (2024) utilizes an LLM itself to generate natural language explanations of its internal representations. It consists of two forward passes, one using a source prompt and the other using an inspection prompt, with a patching operation between them. An inspection prompt is designed to guide an LLM as a probe to extract specific knowledge encoded in its internal representations, aligned with a predefined objective, such as next-token prediction, or attribute extraction. Utilizing an LLM itself with an inspection prompt as a probing mechanism addresses key limitations of prior methods Hernandez et al. (2024); Geva et al. (2022); Belrose et al. (2025), which often rely on predefined probing classes or suffer from limited interpretability due to sub-word tokenization.
In the context of cultural knowledge, these limitations are particularly pronounced. It is challenging to exhaustively define all relevant cultural knowledge. Additionally, a direct projection onto the vocabulary space via the unembedding matrix becomes difficult to interpret since cultural knowledge is frequently tokenized into multiple tokens Naous and Xu (2025). To address these challenges, we introduce a Patchscope-based method tailored for probing the cultural knowledge space. To our knowledge, this is the first work to apply interpretability techniques for investigating cultural knowledge in LLMs.
@@ 4 Probing Cultural Knowledge within Internal Layers
Probing the cultural knowledge processed by each layer for the given input provides insights into how cultural knowledge for one culture is overlapping with different cultures within the inner layers of an LLM. To translate internal representations of LLMs to natural language that reveals the cultural knowledge space, we propose Culturescope, building upon the existing interpretability method, Patchscope Ghandeharioun et al. (2024).
@@ 4.1 Culturescope
Culturescope consists of three stages: inference, scoping-in, and filtering. Our Culturescope allows us to move beyond what is observable from model responses alone, overcoming the limitation of extrinsic evaluation.
@@
An LLM first encodes a tokenized input $T_{i}$ of $i$-th instance and generates an output $O_{i}$, which is an LLM answer to an Open-Ended cultural QA. Then, we compute the representative hidden representations used to generate an answer, which are patched onto the inspection prompt during the scoping-in stage.
As Patchscope does not consider patching with multi-tokens Ghandeharioun et al. (2024), we adopt Bronzini et al. (2024)’s approach originally developed for fact-checking claims to condense an LLM’s cultural answer involving multiple tokens into a single hidden representation. Similarly, we compute the representative hidden representation $x_{*}^{l}$ of the $O_{i}$, which consists of multiple tokens, on the layer $l$. Specifically, we perform the weighted sum of hidden states as in Eq. 2 for the layer $l$. We set the weight $w_{p}$ of each token to one if it is a noun or a verb. Other token weights are set to zero. We use this $x_{*}^{l}$ in the next stage.
$x_{*}^{l}=\sum_{p=0}^{P}x_{p}^{l}*w_{p}$ (2)
@@
At this stage, we further elicit the cultural knowledge encoded in $x_{*}^{l}$ to reveal the cultural knowledge space utilized for the $O_{i}$. To elicit a list of cultural knowledge $\mathcal{CK}_{i}=[ck_{i,1},...,ck_{i,j},$...] from the cultural knowledge space, we design an inspection prompt to elicit cultural knowledge by computing a forward pass. The inspection prompt ends with a placeholder token ‘x’, where we perform the patching as in the Patchscope Ghandeharioun et al. (2024). At the $l$-th layer, we replace the hidden representation at the placeholder token position with $x_{*}^{l}$. The inspection prompt can be found in Appendix B.
@@
To filter out the knowledge that is unrelated to cultural knowledge, we calculate the semantic similarity as an activation score between the input text $T$ and the generated cultural knowledge $ck_{i,j}$. We take a DeBERTA natural language inference (NLI) model to obtain the hidden representation. [^2] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 In Eq. 3, we compute the representative representation of the input text $T$ using the final hidden states from the NLI model. The activation score $s_{i,j}$ of $ck_{i,j}$ is calculated by their cosine similarity.
$g_{t}^{*}=\frac{1}{T}\sum_{t=0}^{T}g_{t}$ (3)
We keep $ck_{i,j}$ when its $s_{i,j}$ is higher than the threshold, which is set to 0.3.

### v2

@@ Abstract
The growing deployment of large language models (LLMs) across diverse cultural contexts necessitates a deeper understanding of LLMs’ representations of different cultures. Prior work has focused on evaluating the cultural awareness of LLMs by only examining the text they generate. This approach overlooks the internal sources of cultural misrepresentation within the models themselves. To bridge this gap, we propose Culturescope, the first mechanistic interpretability-based method that probes the internal representations of different cultural knowledge in LLMs. We also introduce a cultural flattening score as a measure of the intrinsic cultural biases of the decoded knowledge from Culturescope. Additionally, we study how LLMs internalize cultural biases, which allows us to trace how cultural biases such as Western-dominance bias and cultural flattening emerge within LLMs. We find that low-resource cultures are less susceptible to cultural biases, likely due to the model’s limited parametric knowledge. Our work provides a foundation for future research on mitigating cultural biases and enhancing LLMs’ cultural understanding.
@@ 1 Introduction
Social scientists have long observed that globalization often comes at a cost to cultural diversity. This phenomenon, known as ‘Cultural Homogenization,’ occurs when dominant cultures standardize and overgeneralize others, effectively erasing the distinctiveness between cultures Pieterse (1996); Amin (2024). Unfortunately, this imbalance is also reflected in the pre-training data of LLMs, which is predominantly Western-centric (Santurkar et al., 2023). While LLMs do acquire some cross-cultural knowledge during training (Hershcovich et al., 2022; Arora et al., 2023), the imbalance of the sources of cultural knowledge often leads to cultural bias. This results in cultures being either marginalized through underrepresentation or ‘flattened’, wherein their unique traits are conflated with those of dominant cultures. (Nguyen et al., 2023a). For example, when asked about a popular leisure activity for retired men in Azerbaijan, GPT-4 responded with chess, which is a plausible answer, but one that reflects broad post-Soviet stereotypes rather than culturally specific knowledge about Azerbaijan (Myung et al., 2024).
@@
Figure 1: Given the question (about the popular domestic vacation spot in Greece), Culturescope first generates an answer to the cultural question at the Inference stage. Then, it reads the hidden representation from the Inference stage and elicits the cultural knowledge used for the answer (‘Zakynthos’) at the Scoping-in stage. We finalize a list of cultural knowledge after the Filtering stage. Culturescope unveils the internal mechanism of LLMs that cannot be revealed through the Inference stage alone.
These patterns of overgeneralization contribute to what we term ‘cultural flattening’: the model’s tendency to conflate distinct cultural characteristics into dominant cultures. While evaluating model outputs, which corresponds to the extrinsic evaluation, is vital for detecting the presence of such overgeneralization, this approach alone cannot reveal the underlying mechanisms that produce it. It shows us that a model is biased, but not how this bias is encoded in its parameters.
@@
To facilitate the examination of the underlying mechanisms, we propose to study LLMs’ cultural biases with mechanistic interpretability (MI) techniques. MI techniques provide us with methods that can directly examine how cultural biases discovered by the extrinsic evaluation Santurkar et al. (2023) is internally processed within model representations, revealing where and how harmful generalizations emerge. We are the first to propose an approach for intrinsic cultural bias evaluation.
@@
In this work, we introduce Culturescope, a method to probe internal representations and surface the cultural biases in play (Sec. 4.1). Figure 1 illustrates an overview of Culturescope. To examine the intrinsic cultural flattening embedded in the parameter space, we introduce a cultural flattening (CF) score, which quantifies the degree of intersection between cultural knowledge decoded by Culturescope (Sec. 4.2). We implement our framework on two cultural understanding tasks, cultural commonsense Question Answering (QA) and extractive QA, across three different LLMs.
@@
We further challenge the model’s cultural understanding by creating multiple-choice questions (MCQs) with hard negatives (Sec. 3.2.3). Culturally nuanced answers from high-resource cultures or geographically proximate countries are selected as hard negative options to simulate the cultural biases. This setup prevents LLMs from relying on surface-level elimination strategies based on the overgeneralization Khan et al. (2025). Analyzing selected options by LLMs with the attention map method Yuksekgonul et al. (2024) allows us to examine whether extrinsic and intrinsic cultural biases align, by revealing which options the model internally attends to (Sec. 5.1).
Our experiments reveal that LLMs internalize a Western-dominance bias and cultural flattening in their cultural knowledge space, as indicated by asymmetric CF scores between cultures (Sec. 6.1.1) and attention contribution scores (Sec. 6.1.2). This finding is further supported by extrinsic evaluations using MCQs with hard negatives, where LLMs tend to select culturally biased options when answering incorrectly (Sec. 6.2.1). Interestingly, this susceptibility to cultural bias appears weaker for low-resource cultures, likely due to their limited representation in the model’s parametric knowledge. Our findings offer valuable insights for future research aimed at mitigating internalized cultural biases and developing culturally aligned LLMs.
@@ 2 Related Work
Previous work has proposed evaluation datasets and frameworks to assess LLMs’ cultural understanding ability acquired during pre-training Keleg and Magdy (2023); Naous and Xu (2025); Pawar et al. (2025). BLEnD Myung et al. (2024) provides a multilingual commonsense QA dataset spanning 16 countries and regions, designed to uncover cross-cultural disparities in everyday knowledge. CAMeL Naous et al. (2024) compares LLM behavior in Arabic versus Western settings across tasks like story generation, NER, and sentiment analysis, exposing systematic cultural biases in LLMs. Other multilingual benchmarks Zhou et al. (2025); Hasan et al. (2025); Wang et al. (2024a); Cao et al. (2024) construct culturally localized evaluation datasets that span domains such as cuisine, proverbs, news, and reasoning. Across these datasets, performance gaps are consistently observed between high-resource and underrepresented languages and cultures, often linked to pre-training data imbalances that favor dominant regions Naous and Xu (2025).
@@
MI techniques are developed to explain the inner workings of LLMs by identifying responsible model components, such as neurons and attention heads Meng et al. (2022); Geva et al. (2023); Yu et al. (2024). Leveraging their transparency, recent studies have employed MI techniques to investigate how specific behaviors emerge in LLMs. For instance, they have been used to uncover and manipulate components associated with social biases, enabling both diagnostic and steering interventions Liu et al. (2024); Durmus et al. (2024); Yang et al. (2024). Despite growing interest in the cultural capabilities of LLMs, no prior work has explored cultural biases through the lens of MI. Our study fills this gap by applying MI techniques to probe the internal representation of cultural knowledge in LLMs, offering new insights into how cultural understanding is encoded and organized within the model.
@@ 3.1 Preliminaries
For each culture of interest $y$, a dataset $D=\{(q_{0},C_{0}),...,(q_{N-1},C_{N-1})\}$ consists of $N$ tuple instances that contain a question $q$ and an option list $C=[c_{0},c_{1},c_{2},a]$ of one gold answer $a$ and three other options for MCQ. For MCQ, an LLM is given $C$ and $q$ to generate an output $O=[o_{0},...,o_{P-1}]$ consisting of $P$ tokens. For open-ended QA, an LLM is only given $q$ to generate the output $O$.
To generate an answer, an LLM converts a tokenized input text $T=[t_{0},...,t_{S-1}]$ containing $S$ number of tokens into $d$-dimenstional vectors using the embedding matrix $E\in\mathbb{R}^{|\mathbb{V}\times d|}$, where $\mathbb{V}$ denotes the number of vocabularies. Then, the vectors are processed through $L$ layers, each containing a multi-head self-attention (MHSA) layer and an MLP layer. The hidden representation $x_{i}^{l}$ from a layer $l$, on a token $t_{i}$ is computed by:
@@ 3.2.1 Cultural QA Datasets
We select BLEnD Myung et al. (2024), a cultural commonsense QA dataset, and CAMeL-2 Naous and Xu (2025), an extractive QA dataset featuring culturally grounded entities. BLEnD Myung et al. (2024) is a hand-crafted multilingual benchmark designed to evaluate LLMs’ everyday knowledge across diverse cultures. CAMeL-2 Naous and Xu (2025) is a bilingual benchmark originally constructed to evaluate LLMs’ entity extraction capabilities on Arabic and English entities. We use English and Spanish questions from BLEnD and use both Arabic and English from CAMeL-2. For both datasets, we select 14 cultural groups to keep a similar culture distribution between two datasets. Dataset details can be found in Appendix A.
@@ 3.2.2 Grouping of Cultures
We categorize 14 countries from each dataset along the resource dimension and the region dimension to study how overgeneralization manifests across these dimensions. For the resource dimension, we adopt the taxonomy proposed by Joshi et al. (2020), which we simplify into three resource levels. For the region dimension, we group countries into six regions based on continents. For more details, we refer to Appendix A.1.
@@ 3.2.3 Cultural MCQ with hard negatives
Khan et al. (2025) found that if MCQs lack the adversarial depth to probe genuine cultural understanding, models can exploit surface-level elimination strategies without understanding cultural distinctions. Thus, we propose a cultural MCQ with hard negatives to study how overgeneralization—driven by regional or resource dominance or similarity—affects the downstream task. Since BLEnD Myung et al. (2024) provides different answers from each culture with the same question, we create BLEnD-resource and BLEnD-region partition using culturally nuanced answers in BLEnD.
@@
For dataset creation, we design two types of multiple choice question option lists that incorporate hard negative options: $C_{resource}$ and $C_{region}$, corresponding to BLEnD-resource and BLEnD-region, respectively. For $C_{resource}$, given a question $q$ targeting culture $y$, we sample one culture from each of the three resource levels excluding $y$. We obtain these three cultures’ respective gold answers when substituted into $q$ for $y$, resulting in three hard negative options: $c_{high}$, $c_{mid}$, and $c_{low}$. For $C_{region}$, we sample one culture from the same geographical region as $y$ (excluding $y$) and extract its corresponding answer to construct a region-based hard negative option, $c_{sameregion}$. Two additional options ($c_{region1}$, $c_{region2}$) are randomly selected from countries in different regions. We report descriptive statistics of the number of selected countries to demonstrate the balanced sampling in Appendix A.2. All options are shuffled to avoid positional bias.
Figure 2 shows the example of $C_{resource}$ and $C_{region}$ for the question about China’s popular family game. Cultural MCQ with hard negative options allows us to examine when a model generates incorrect answers to cultural questions, whether a model’s cultural confusion arises from similarity in resource level or regional proximity.
@@ 3.3 Models
Application of MI methods requires full access to model weights, hence we conduct our experiments with three recent open-sourced LLMs: Meta-Llama-3.1-8B-Instruct (Llama-3.1, Grattafiori et al. (2024)), aya-expanse-8b (aya-expanse, Dang et al. (2024)), and Qwen2.5-7B-Instruct (Qwen2.5, Qwen Team (2024)).
@@ 3.4 Patchscope
Patchscope Ghandeharioun et al. (2024) utilizes an LLM to generate natural language explanations of its internal representations via two forward passes with a patching operation in between. An inspection prompt serves as a probe to extract specific knowledge encoded in the model’s internal states, aligned with a predefined objective, such as next-token prediction, or attribute extraction. Utilizing an LLM itself with an inspection prompt as a probing mechanism addresses key limitations of prior methods Hernandez et al. (2024); Geva et al. (2022); Belrose et al. (2025), which often rely on predefined probing classes or suffer from limited interpretability due to sub-word tokenization.
These limitations are particulary pronounced for cultural knowledge, which is difficult to enumerate exhaustively and often spans multiple tokens Naous and Xu (2025). To address these challenges, we introduce a Patchscope-based method tailored for probing the cultural knowledge space. To our knowledge, this is the first work to apply interpretability techniques for investigating cultural knowledge in LLMs.
@@ 4 Probing Cultural Knowledge within Internal Layers
Probing the cultural knowledge processed by each layer for the given input provides insights into how cultural knowledge for one culture is overlapping with different cultures within the inner layers of an LLM. To translate internal representations of LLMs to natural language that reveals the cultural knowledge space, we propose Culturescope, building upon the existing interpretability method, Patchscope Ghandeharioun et al. (2024). Culturescope consists of three stages: inference, scoping-in, and filtering. Culturescope allows us to move beyond what is observable from model responses alone, overcoming the limitation of extrinsic evaluation.
@@ 4.1 Culturescope
An LLM first encodes a tokenized input $T_{i}$ of $i$-th instance and generates an output $O_{i}$, which is an LLM answer to an open-ended cultural QA consisting of $P$ number of tokens. However, activation patching during the scoping-in stage requires a single-token representation of the $O_{i}$.
Since Patchscope does not consider patching with multi-tokens Ghandeharioun et al. (2024), we adopt Bronzini et al. (2024)’s approach, originally developed for fact-checking claims, to condense an LLM’s cultural answer involving multiple tokens into a single hidden representation. Specifically, to compute the representative hidden representation $x_{*}^{l}$, we perform the weighted sum of hidden states as in Eq. 2 for the layer $l$. We set the weight $w_{p}$ of each token to one if it is a noun or a verb. Other token weights are set to zero. The resulting $x_{*}^{l}$ is then patched onto the inspection prompt during the scoping-in stage.
$x_{*}^{l}=\sum_{p=0}^{P-1}x_{p}^{l}*w_{p}$ (2)
@@
During this stage, we utilize activation patching to elicit the cultural knowledge encoded in $x_{*}^{l}$, revealing the cultural knowledge space utilized for generating $O_{i}$. To do this, the LLM performs another inference using an inspection prompt, which generates a comma-separated sequence of cultural knowledge items. The inspection prompt ends with a placeholder token ‘x’, where patching is done following Patchscope Ghandeharioun et al. (2024). At layer $l$, we replace the hidden representation at the placeholder token position with $x_{*}^{l}$. The inspection prompt can be found in Appendix B. Finally, we split the generated sequence into a list of cultural knowledge $\mathcal{K}_{i}=[k_{i,1},...,k_{i,j},$...].
@@
To filter out the knowledge that is unrelated to cultural knowledge, we calculate the semantic similarity as an activation score between the input text $T$ and the generated cultural knowledge $k_{i,j}$. We take separate encoder-based semantic search models for each language to obtain the hidden representation. In Eq. 3, we compute the representation of the input text $H_{T}=[h_{0},..h_{s},..h_{S-1}]$ using the final hidden states from the semantic search model. The activation score $e_{i,j}$ of $k_{i,j}$ is calculated by the cosine similarity.
$h_{t}^{*}=\frac{1}{S}\sum_{s=0}^{S-1}h_{t}$ (3)
We keep $k_{i,j}$ when its $e_{i,j}$ is higher than the average of all $e$. Details for implementation is provided in Appendix C.